a 2.0 m length of string with a mass density of 2.95 x 10^-4 kg/m is fixed at both ends and driven at 120 Hz. The tension is varied to obtain standing waves (resonance) on the string.
1. what is the longest wavelength for a standing wave possible on the string?
2. the tension on the string is varies to obtain fourth harmonic
a. what is the wavelength of this standing wave?
b. what is the wave speed
3. what should the tension be to obtain the third harmonic?

a 2.0 m length of string with a mass density of 2.
Answer #1

1. Longest wavelength is 4m as the longest wave has a single antinide.
2. a. Wavelength is 2m/5 =0.4 m
   b. v = frequency* wavelength = 48 m/s
3. T = v^2 * mass density = 0.68 N